IT Service Outage History

UPDATE: Partial Cable TV Outage

Submitted on Monday, 9/8/2014, at 9:18 AM

Satellite dishes will be moved at about 12:00 noon today to make way for the greenway dormitory project.  The outage will not affect all channels, but it will include TV Japan, local channels and all HD channels.  Once the channels disappear from the line-up, they will be off-line for approximately two hours.

2:30PM UPDATE:  The outage was delayed until 2:00PM.  HD channels are off-line now, but should be restored soon.  Then TV Japan and the Arabic Channel will go off-line for a period lasting from one to two hours.

MONDAY, SEPTEMBER 8 UPDATE:  The cable TV lineup was restored by 4:30PM on Friday, but the dish relocation work was not completed, due to an out-of-box equipment failure.  Contractors are expected to return on Tuesday, September 9.  A Tuesday outage is expected to impact TV Japan, the Arabic Channel and possibly the local channels.

386682
Unknown
Greenway Project Preparation
Friday, September 5, 2014, at 5:00 PM
Friday, September 5, 2014, at 2:00 PM to at 5:00 PM

ACData

Submitted on Wednesday, 9/3/2014, at 12:59 PM

ACData was unexpectedly unavailable and required a server reboot.  

387590
Unknown
A damaged file needed to be repaired.
Wednesday, September 3, 2014, at 11:00 AM to at 12:45 PM

Files Server Maintenance, Thursday, August 21st.

Submitted on Friday, 8/22/2014, at 4:10 PM

As part of ongoing preparations for the start of the new semester, the main file servers (FS001 and FS005, serving the main U: and I: drives on Windows, Userfiles on Macs, as well as the STORAGE server) will be taken down for maintenance tomorrow evening (Thursday, 8/21) from 9:00 PM to 11:00 PM.

Users are advised not to have any documents on these servers open during this period, and to close any documents before leaving at the end of the day on Thursday.

All systems should be back up by 11:00 PM on Thursday night.

Note that this outage is only for the file servers -- all other service including E-mail, Moodle, and the CMS will continue to run normally.

386713
386714
386715
Unknown
system maintenance
Thursday, August 21, 2014, at 11:00 PM
Thursday, August 21, 2014, at 9:00 PM to at 11:00 PM

COMPLETE: Scheduled outage for PON software upgrade

Submitted on Wednesday, 7/30/2014, at 6:32 PM

UPDATE:  This upgrade was completed successfully and the vulnerability has been eliminated.

 

A scheduled outage will affect services offered by Amherst College Information Technology.

On Wednesday, July 30, at 5:00 PM, there will be a loss of network access and phone service at locations served by our passive optical network (PON), while we perform a critical software upgrade to fix the bug responsible for the unscheduled outages on Monday, July 28.  Affected areas include: Johnson Chapel, College Hall, 79 South Pleasant Street, Pratt Field faciltities building, Whalen House, parts of the gym, parts of Keefe Campus Center, parts of the Converse basement, Frost Barker Room, and the new grounds building. 

We appreciate your patience while we address this vulnerability.

-- Stephen A. Judycki   sajudycki@amherst.edu  413-542-5823

    Director, IT Communications Infrastructure, Amherst College

387583
386672
386677
386676
386678
386674
387603
386675
Unknown
Wednesday, July 30, 2014, at 6:00 PM
Wednesday, July 30, 2014, at 5:00 PM to at 6:15 PM

RESOLVED: File Server FS001 Back Online

Submitted on Wednesday, 7/30/2014, at 10:54 AM

An ongoing outage is affecting services offered by Amherst College Information Technology. We are working on the problem and appreciate your patience.

FS001 is down. It contains the "I: drives" of many administrative departments and the "U: drives" of many administrative staffers. Shares on the server will be unavailable until service is restored.

Update: FS001 was rebooted about 9:15 this morning and is working normally. Please call the Help Desk if you have any further issues with the server.

386713
386714
Unknown
Tuesday, July 29, 2014, at 8:00 AM to at 9:15 AM

Resolved: Morgan Hall Network Outage

Submitted on Tuesday, 7/29/2014, at 9:09 AM

UPDATE: Defective network switch replaced.  We will continue to monitor, but we believe the problem has been resolved.

An ongoing outage is affecting services offered by Amherst College Information Technology. We are working on the problem and appreciate your patience.

Weeklong intermittent problems with the network in Morgan Hall became more pronounced today.  The network is hard down and we are working on a fix.  We apologize for the inconvenience and disruption.

386677
386676
386678
386674
386675
Unknown
Monday, July 28, 2014, at 9:30 AM to Tuesday, July 29, 2014, at 8:45 AM

RESOLVED: PON Network Problems

Submitted on Monday, 7/28/2014, at 5:25 PM

UPDATE:  The problem that impacted network and telephone service today at Johnson Chapel, College Hall, 79 South Pleasant, and parts of other buildings such as Keefe Campus Center, has been resolved.  It involved a strange set of circumstances that included a software bug that was triggered by someone plugging in a telephone incorrectly.  What ensued was a traffic storm that prevented normal communications from taking place.  The vulnerability (the bug) will exist until an after-hours software upgrade can be scheduled, perhaps one day next week.  We apologize for the length of time it took to correct this problem.

 

An ongoing outage is affecting services offered by Amherst College Information Technology. We are working on the problem and appreciate your patience.

We have been experiening sporadic service issues with the PON network throughout the afternoon.  We believe the problem has been discovered.  There will be an unavoidable hard outage of the entire PON network at 4:30PM lasting for approximately 10 minutes.  We apologize for the disruptions.

Areas affected include the Johson Chapel, College Hall and 79 South Pleasant buildings and some offices in Keefe Campus Center.

 

386672
386677
386676
386678
386674
386675
Unknown
Monday, July 28, 2014, at 11:00 AM to at 5:00 PM

RESOLVED: Morgan Hall Network Outage

Submitted on Wednesday, 7/23/2014, at 4:09 PM

Update:  fiber cleaning was completed by 9:00 AM and logs have shown the network to be error-free since that time.

The network was restored at 8:15 AM, but error logs indicate that there may be an underlying cable/media problem.  Technicians will be cleaning fiber optic connections at several points along the path this morning, which will cause the network in Morgan Hall to go down and up at least several times, until this work has been completed.  We apologize for the disruption and/or invconvenience.

386677
386676
386678
386674
387603
386675
Unknown
Wednesday, July 23, 2014, at 7:30 AM to at 9:00 AM

RESOLVED: Morgan Hall Network Outage

Submitted on Monday, 7/21/2014, at 6:13 PM

Network connectivity has been restored to the building, but there may be lingering wireless problems that will be addressed on 7/22/14.

Unknown
Monday, July 21, 2014, at 3:39 PM to at 6:00 PM

ACDATA offline

Submitted on Friday, 7/18/2014, at 9:05 AM

The ACDATA system is currently offline. We are troubleshooting the problem.

 

Service restored.

Unknown
Friday, July 18, 2014, at 8:14 AM to at 8:20 AM